### v1

# Science and Technology

Abstract
We study zero-shot instance navigation, in which the agent navigates to a specific object without using object annotations for training. Previous object navigation approaches apply the image-goal navigation (ImageNav) task (go to the location of an image) for pretraining, and transfer the agent to achieve object goals using a vision-language model. However, these approaches lead to issues of semantic neglect, where the model fails to learn meaningful semantic alignments. In this paper, we propose a Prioritized Semantic Learning (PSL) method to improve the semantic understanding ability of navigation agents. Specifically, a semantic-enhanced PSL agent is proposed and a prioritized semantic training strategy is introduced to select goal images that exhibit clear semantic supervision and relax the reward function from strict exact view matching. At inference time, a semantic expansion inference scheme is designed to preserve the same granularity level of the goal-semantic as training. Furthermore, for the popular HM3D environment, we present an Instance Navigation (InstanceNav) task that requires going to a specific object instance with detailed descriptions, as opposed to the Object Navigation (ObjectNav) task where the goal is defined merely by the object category. Our PSL agent outperforms the previous state-of-the-art by 66% on zero-shot ObjectNav in terms of success rate and is also superior on the new InstanceNav task. Code will be released at https://anonymous.4open.science/r/PSL/.
Keywords:
Zero-Shot Instance Navigation Zero-Shot Object Navigation Image Goal Navigation Prioritized Semantic Learning
1 Introduction
Visual navigation is a fundamental skill for embodied agents to travel efficiently in complex 3D environments [65]. Among different kinds of visual navigation tasks, Zero-shot Object Navigation (ZSON) is a promising task and paves the way for developing a general embodied agent [2, 54, 49, 12, 60, 13, 27, 47, 24], since it requires zero scene object annotations for training. However, the criteria of the current ZSON task is to evaluate whether the agent can go to the object that matches solely the given category, which remains a considerable gap from real-life applications that require the identification of a specific object. Therefore, we propose to extend ZSON to Zero-shot Instance Navigation (ZSIN) task in the HM3D environment [42, 58], which requires the agent to go to a unique object instance given detailed text descriptions. Both the ZSON and ZSIN tasks are very challenging due to the zero-shot constraint, where the object goal, required by the reward function in standard reinforcement learning frameworks such as PPO [45, 53], can not be trivially defined.
Figure 1: Pilot studies on four navigation agents: Layout-Only (LO) agent, Semantic-Only (SO) agent, ZSON [37] agent, and our PSL agent. We intuitively think that the agent with a fixed semantic observation encoder i.e., CLIP, possesses good semantic perception ability, while a learnable observation encoder delivers great navigation capacity. (a) The agent trained on ImageNav task does not necessarily need to learn the semantic information to obtain a high success rate. (b)-(c) Our PSL agent achieves both strong semantic perception and navigation capacity. [^2] The LO agent does not support InstanceNav task due to the lack of a semantic goal encoder.
To achieve zero-shot visual navigation, pioneering works [1, 37] turn to pre-trained vision-language models such as CLIP [41] for its zero-shot capability and propose to train the agent on Image-goal Navigation (ImageNav) pretext task that requires the agent to go to a randomly sampled goal image [65]. To be specific, using the CLIP vision encoder to obtain a semantic goal from the goal image for training, the ZSON agent [37] can be transferred to ObjectNav by switching the semantic goal into the CLIP text embedding. Although the ZSON agent establishes a solid semantic agent with good navigation capacity, we argue that the above pipeline does not deliver a navigation agent with strong semantic perception ability.
[…]
In this paper, we propose a Prioritized Semantic Learning (PSL) method to improve the semantic perception and understanding ability of the navigation agent. Our PSL approach is comprised of three parts: a PSL agent architecture, a prioritized semantic training strategy, and a semantic expansion inference scheme. First, we equip the agent with an additional CLIP vision encoder to encodes the semantics in the observation. Subsequently, a Semantic Perception Module (SPM) is proposed to comprehend the semantic differences between the observation and goal images. Second, we take advantage of the entropy-minimization technique to select goal images with clear semantic supervision. Moreover, we relax the reward function to focus more on semantic correspondence rather than strict geometric matching Third, considering that the proposed SPM learns to comprehend the semantic differences of image embeddings during training, we develop a semantic expansion inference scheme that retrieves image embeddings using text queries. The retrieved image embeddings expand the text embeddings with rich visual priors, facilitating more precise indications. Without bells and whistles, the proposed PSL method achieves state-of-the-art performance on both widely-evaluated ObjectNav and our extended InstanceNav benchmarks. For example, in the ObjectNav task, PSL outperforms the ZSON baseline [37] and the ESC method [63] that uses large-language models by 16.9% and 3.2% success rate, respectively. Furthermore, we verify the proposed PSL agent has a stronger semantic perception ability by visualizing the semantic similarities between the agent’s last observation and the goal images in Fig. 2 (c).
Our main contributions are summarized as follows:
  • We investigate the semantic perception ability of different navigation agent architectures and reveal that the commonly-used ImageNav pre-training task does not necessarily require the agent to learn semantic information.
  • We propose a prioritized semantic learning method that includes a strong agent architecture, a novel training strategy and an effective inference scheme to improve the agents’ semantic perception ability in zero-shot object/instance navigation tasks.
  • Extensive experiments and sufficient ablation studies on both ObjectNav and InstanceNav benchmarks demonstrate the superior performance of the proposed PSL method over the state-of-the-art.
2 Related Work
2.1 Visual Navigation
Visual navigation is a fundamental and crucial task for robots to perform various embodied tasks [25, 4, 52, 46, 26]. In visual navigation tasks, the agent receives a geometry [44, 53] or semantic [5, 65] goal and an RGB observation from the camera attached to the agent, and then makes efforts to navigate to this goal according to the observation in each time step. These tasks can be divided into different categories, including point-goal navigation [44, 53], object-goal navigation [57, 56, 37, 63, 6, 43], image-goal navigation [65, 57, 56, 1, 10, 15, 39, 23, 18, 55], and vision-language navigation [3, 22, 21, 19, 8], etc.  Taking image-goal navigation [55] as an example, the agent needs to explore the environment, locate the target object specified by the image, and navigate to its proximity. A common trail for these visual navigation tasks is to train the agent in the simulator [44, 48] with photo-realistic environments [42, 58, 7] and physical engine, and provide the agent with the goal indication as well as its corresponding trajectory annotation. Since labeling objects in the 3D scenes is labor expensive, recent works [1, 37, 63, 6] have explored how to design an object-goal navigation agent without human annotations. These methods are easy to scale up [37] and show potential to solve the open set challenge [36] in locating objects. In this work, we focus on a more challenging task of visual navigation, zero-shot instance navigation, in which the agent is required to navigate to an object specified by detailed descriptions, for example, an instance image [20] or text.
2.2 Vision-Language Model
Vision-language models (VLM) provide mutual understanding for both vision and language modalities [50] and thereby naturally serve as a fundamental bedrock for vision navigation. To build a visual navigation agent, two lines of work are widely employed, namely, representation-based and module-based methods. Representation-based methods [41, 17, 29] aim to co-embed the image and text into a common semantic space. CLIP [41] and ALIGN [17] apply contrastive loss to train the image and text encoders on large-scale pair-wise image text datasets. ALBEF [29] further proposes to leverage a momentum distillation mechanism that alleviates the weak correlation between noisy image text pairs to improve multimodal semantic learning. As for the module-based methods [30, 28, 62, 14, 34, 33], they aim to learn a strong multimodal reasoner to simultaneously comprehend vision-language modalities and produce text responses, which are commonly used as the navigation controller [63, 6]. Early works such as VisualBERT [30] apply mask language loss on both image patches and word embeddings to pre-train a multimodal reasoning model. To incorporate the strong reasoning ability of large language models, recent works like BLIP-2 [28], LLaMA-Adapter [62, 14] and LLaVA [34, 33] perform visual instruction tuning on pretrained large language models. However, using these module-based vision language models is compute-intensive, making them unsuitable for real-life robotic scenarios that require nearly real-time response. In this paper, we construct the navigation agent using the representation-based VLM and aim to improve the semantic perception ability of the agent.
3 Motivation of Prioritized Semantic Learning
ImageNav is an important pre-training task to achieve Zero-Shot Instance Navigation (ZSIN) as it provides supervision signals to navigate in 3D environments with visual observation and a semantic indication [37]. However, we argue that the ImageNav task can be solved without semantic clues, resulting in a sub-optimal pre-training objective. To verify this, we conduct a preliminary experiment on the HM3D ImageNav dataset that is used to train a zero-shot object navigation agent in previous literature [37]. We set up four different agents, namely the Layout-Only (LO) agent, Semantic-Only (SO) agent, ZSON agent, and our PSL agent. Among them, the Semantic-Only agent is equipped with two frozen CLIP ResNet-50 encoders to obtain semantic-level observations and goals. The Layout-Only agent is adopted with two trainable ResNet-50 encoders and we remove the semantic information in both observation and goal images by using a Sobel operator to calculate gradient images, which leave only layout information like contours and edges (see Fig. 2 for more details). Results in Fig. 2 (a) show that the agent does not necessarily need to learn semantic information to arrive at an image goal, merely relying on the layout information to perform view matching can obtain a high success rate.
4 Prioritized Semantic Learning
4.1 Preliminaries
We follow ZSON [37] to adopt image-goal navigation pre-training and perform the zero-shot evaluation on downstream tasks. During pre-training, the image-goal navigation (ImageNav) objective provides supervision signals to update the agent. Specifically, for each episode with a randomly sampled start point and goal point, the agent is asked to navigate to the goal destination according to a goal image. Empowered by the cross-modal alignment ability of Vision-and-Language models (i.e., CLIP [41]), we can evaluate the trained agent on the object-goal navigation (ObjectNav) dataset by replacing the image embeddings with text embeddings.
4.2 Prioritized Semantic Learning Agent Architecture
We propose a Prioritized Semantic Learning (PSL) agent that emphasizes the reasoning of the semantics differences in the goal and observation. The key components of PSL agent include observation and goal encoders, a semantic perception module module and a recurrent policy network.
Figure 2: Overview of our PSL. 1) During ImageNav pre-training, we provide clear semantic supervision signals with the Prioritized Semantic Training strategy. 2) Our PSL agent exploits the Semantic Perception Module (SPM) to achieve both strong semantic understanding and navigation capacity. 3) During inference, a Semantic Expansion Inference scheme is incorporated to ensure the same semantic granularity of the goal-embedding between training and testing.
Observation and Goal Encoders.
Following the ZSON baseline, given the observation image ${\bf I}_{\mathrm{O}}$ and the goal image ${\bf I}_{\mathrm{G}}$, we apply a learnable ResNet50 encoder and a fixed CLIP encoder to obtain the observation embedding ${\bf z}_{\mathrm{O}}$ and the goal embedding ${\bf z}_{\mathrm{G}}$, respectively. Since with only a trainable ResNet50 as the observation encoder may not effectively learn the semantic information, we additionally leverage a frozen CLIP encoder to extract semantic-level observation ${\bf z}_{\mathrm{S}}$. Note that the ResNet50 are initialized from the one pre-trained by OVRL [57].
Semantic Perception Module.
We then introduce the semantic perception module, an MLP layer that encodes the semantic correspondence between ${\bf z}_{\mathrm{G}}$ and ${\bf z}_{\mathrm{S}}$ into a low-dimension feature. Specifically, it takes both ${\bf z}_{\mathrm{G}}\in\mathbb{R}^{C_{1}}$ and ${\bf z}_{\mathrm{S}}\in\mathbb{R}^{C_{1}}$ as input and produce semantic perception embedding ${\bf z}_{\mathrm{SP}}\in\mathbb{R}^{C_{2}}$, where $C_{2}<2\times C_{1}$. The semantic perception module reduces the feature dimension, serving as a bottleneck to condense the critical semantic correspondence in both the goal image and observation image.
Policy Network.
Based on the semantic perception embedding ${\bf z}_{\mathrm{SP}}$ and the observation embedding ${\bf z}_{\mathrm{O}}$, we train a navigation policy $\pi_{\theta}$ using reinforcement learning (RL):
${\bf s}_{t},{\bf h}_{t}=\pi_{\theta}({\bf z}_{\mathrm{SP}}\oplus{\bf z}_{\mathrm{O}}\oplus{\bf a}_{t-1}|{\bf h}_{t-1})$ (1)
where ${\bf s}_{t}$ is the embedding of the agent’s current state, which is then used to predict an action distribution among six action categories, including MOVE_FORWARD, TURN_LEFT, TURN_RIGHT, STOP, LOOK_UP and LOOK_DOWN. ${\bf h}_{t-1}$ denotes hidden state of the recurrent layers in policy $\pi_{\theta}$ from the previous step. Following previous methods [1, 37], we adopt an actor-critic network to predict state value $c_{t}$ and action $a_{t}$ using $s_{t}$ and train it end-to-end using PPO [45]. More details can be found in the Appendix.
Figure 3: Illustration of the perspective relaxation and goal view selection approach in our training strategy. The goal view image set is expanded with additional yaw and tilt views using perspective relaxation. Top-10 views with minimum entropy are circled in red with their ranking.
4.3 Prioritized Semantic Training Strategy
Previous literature has pointed out that a large portion of image goals in the ImageNav task lead to ambiguity (e.g. looking at walls). We provide statistical analysis on the
Figure 4: Category distribution of goal images in the ImageNav task.
distribution of ImageNav goal images to support this speculation. In particular, We select 6 object categories that widely exist in HM3D scenes, as well as two additional categories “wall” and “room” to represent images with ambiguous semantic content. Then we use CLIP to perform zero-shot classification on all goal images by selecting the category with maximum image-text similarity. In Fig. 4, in the original HM3D ImageNav dataset, most of the goal images cannot be categorized as common object categories; instead, they are classified as meaningless indications (i.e., “wall” and “room”). We believe that the unreasonable distribution of goal images aggravates the semantic neglect issue.
Entropy-minimized Goal View Selection.
To alleviate the semantic neglect issue, we propose an entropy-based goal view selection method to ensure a dominant object exists in the goal images. Specifically, we randomly select a base orientation at the goal point and rotate the agent for $\Omega$ times uniformly to render images in different view angles. Then we obtain image embeddings $\{{\bf v}_{\omega}\mid\omega\in\Omega\}$ using CLIP [41] vision encoder and calculate the similarity between image embeddings with all text embeddings of the candidate object categories. We follow the standard setting of Object-Goal navigation approaches [57, 56, 63] to use 6 common objects ${\mathcal{C}}$ in the 3D simulator and compute query vectors $\{{\bf q}_{c}\mid c\in{\mathcal{C}}\}$ using CLIP text embedding. The objective is to select a goal view $\omega^{*}$ that minimizes the entropy of probability distribution among all object categories:
$\omega^{*}=\underset{\omega\in\Omega}{{\arg\min}}-\frac{1}{\log(|{\mathcal{C}}|)}\sum_{c\in{\mathcal{C}}}{\bf p}_{c}{\log{\bf p}_{c}},$ (2)
where ${\bf p}_{c}=\mathrm{softmax}\left(g({\bf v}_{w},{\bf q}_{c})\right)$ is the class probability w.r.t. $c^{\text{th}}$ object category, $g({\bf a},{\bf b})=\tau\cdot\frac{{\bf a}^{T}{\bf b}}{\lVert{\bf a}\rVert_{2}\lVert{\bf b}\rVert_{2}}$ is the scaled cosine similarity function and $\tau$ is the temperature. As shown in Fig. 4, the goal images with ambiguous semantics is significantly reduced. Additionally, image distribution among all object categories becomes more balanced.
Perspective Reward Relaxation.
Choosing goal views with rich semantic content enables the agent to focus on meaningful regions in the goal images. However, the ImageNav reward [1, 37] still motivates the agent to strive for an exact match between the goal image and the observation image. We tackle this challenge by introducing a perspective reward relaxation technique that cancels exact matching. Specifically, we augment the goal view selection process by rendering additional circles of images around the agent at different pitch angles. As illustrated in Fig. 3, we then perform goal view selection among all rendered views that vary in pitch and yaw angles. Furthermore, we rewrite the reward function of PPO training by relaxing the perspective matching requirement:
$\begin{split}R^{\text{PSL}}_{t}=&\underbrace{\gamma^{suc}*\mathds{1}\{d_{t}<\epsilon^{d}\}}_{\textit{reach the goal location or not}}+\underbrace{\gamma^{suc}*\mathds{1}\{d_{t}<\epsilon^{d}\}*\mathds{1}\{(\mathrm{extract}_{\mathbf{Y}}({\bf a}_{t})<\epsilon^{a}\}}_{\textit{match the goal view or not}}\\ &+\underbrace{r_{d}(d_{t},d_{t-1})+\mathds{1}\{d_{t}<\epsilon^{d}\}*\mathrm{extract}_{\mathbf{Y}}(r_{a}({\bf a}_{t},{\bf a}_{t-1}))}_{\textit{closer to the goal or not}}-\gamma^{delay},\end{split}$ (3)
which is comprised of four parts. First, a success reward $\gamma^{suc}=5$ is given when distance to goal metric $d_{t}$ is smaller then threshold $\epsilon^{d}$. Second, an angle success reward is given when the agent reaches the goal location and faces the goal orientation, where $\mathrm{extract}_{\mathbf{Y}}({\bf a}_{t})$ means the component of quaternion ${\bf a}_{t}$ rotating around the $\mathbf{Y}$ axis. We only encourage the agent to head to the target in the x-z plane while ignoring its pitch heading. Third, a dense reward compose of reduced distance $r_{d}$ and angle $r_{a}$ to the target at time step $t$. Last, a delay penalty $\gamma^{delay}$ is to encourage efficiency.
4.4 Semantic Expansion Inference Scheme
In the inference stage, the agent receives a text description of the target object. Existing methods directly replace the image-goal embedding with the text-goal embedding of the input text description to perform zero-shot evaluation. However, the granularity gap [32] between image and text embeddings could considerably impair the evaluation performance [31]. Since it is non-trivial to obtain text annotations for all goal images (28.8M) during training, we opt to bridge the modality gap by augmenting the text queries with semantic-consistent image embeddings from a pre-constructed support set. We name this method semantic expansion inference.
During training, we maintain a support set ${\mathcal{S}}=\{{\bf s}\in\mathbb{R}^{d}\mid{\bf s}_{i}\neq{\bf s}_{j},g({\bf s}_{i},{\bf s}_{j})<\lambda\}$. The threshold is set $\lambda=0.8$ to store representative image embeddings, resulting in around 0.1M vectors. These embeddings approximate the distribution of goal images, and they encoded rich information from a variety of object instances, regarding their detailed attributes (e.g., color, shape, texture, and material). During testing, given a language description of the target object, we leverage the CLIP text encoder to produce a feature ${\bf z}_{\mathrm{T}}$. Then we query a goal embedding using the retrieval operation:
${\bf z}_{\mathrm{R}}=\sum_{{\bf s}_{i}\in{\mathcal{S}}}\frac{\mathrm{exp}(g({\bf z}_{\mathrm{T}},{\bf s}_{i}))}{\sum_{{\bf s}_{j}\in{\mathcal{S}}}\mathrm{exp}(g({\bf z}_{\mathrm{T}},{\bf s}_{j}))}\ast{\bf s}_{i}.$ (4)
We perform a weighted sum based on the similarity score to avoid inferior performance produced by the nearest neighbor querying [31].
5 Experiments
In this section, we first compare the proposed PSL agent with previous zero-shot methods in the ObjectNav task that target finding arbitrary object instances of the same category, and then investigate how it performs when instructed to navigate to a specific object in the InstanceNav task. We also conduct extensive ablation experiments to empirically study the effectiveness of different modules in our method.
5.1 Experimental Setup
Datasets.
During pre-training, we utilize the 7,200,000 episodes generated by ZSON [37] from HM3D [58]. We randomly select 4 goal images for each episode from 10 candidates with minimal entropy in Eqn. 2. For ObjectNav evaluation, we adopt the habitat2022 challenge dataset with 2000 episodes and 6 object categories. For InstanceNav, we first evaluate the agent on IIN dataset [20], an image-goal track with 795 unique instances in 1000 test episodes, and then extend it to the text-goal setting.
Agent Configuration.
We follow the configuration in [37] with an agent height of 0.88m, radius of 0.18m, and a single $640\times 480$ RGB sensor with a ${}^{\circ}$ horizontal field-of-view (HFOV) placed 0.88m from the ground. For InstanceNav, in the image-goal setting, the agent receives a goal image captured in a random camera configuration that does not match the agent’s own. In the text-goal setting, the agent receives text description of both intrinsic and extrinsic attributes (e.g., “Find a double bed and its material is a beige bamboo frame. There are only two objects around the bed: a table and a window."). For ObjectNav, the agent receives the category text as instruction like “Find a bed”.
Evaluation Metrics.
We evaluate the agent’s performance by two metrics: Success Rate (SR) and Success Rate Weighted by Path Length (SPL). An episode is marked as a success if the agent navigates to the goal object and executes the STOP action within 0.1m to any viewpoints of the goal object. The SR metric is the average number of success indicators among all episodes. The SPL metric is defined as $\frac{1}{N}\sum_{i=1}^{N}\textrm{SR}_{i}\frac{l_{i}^{*}}{\mathrm{max}(l_{i},l_{i}^{*})}$, where $l_{i}$ denotes to the agent’s path length and $l_{i}^{*}$ denotes to the ground-truth path length. We report the average number over three runs with different random seeds.
5.2 Experiments on the ObjectNav Task
In Table 1, we report the zero-shot evaluation results on the HM3D ObjectNav task. Our PSL agent outperforms the ZSON baseline [37] by $16.9\%$ in SR ($42.4\%$ vs. $25.5\%$), thanks to our training strategy and inference scheme design. More importantly, our method exceeds the LLM-based methods in the first time. Compared to the PixelNav method [6] that learns to navigate to a point deduced by a powerful multi-modal LLM [40], our PSL agent achieves $+4.5\%$ improvement in SR. Furthermore, taking into account mapping-based methods that perform LLM-guided frontier-based exploration [59], we see $+6.0\%$ improvements against L3MVN [61] and $+3.2\%$ against ESC [63]. Our method shows great potential to be applied in the real environment as it does not require accurate GPS coordinates to build a map. Besides, our LLM-free architecture is more efficient during inference.
5.3 Experiments on the InstanceNav Task
The ObjectNav task assigns multiple available destinations for each episode. It is not sure whether the agent can distinguish object instances. In comparison, we extend the zero-shot object navigation task to the zero-shot instance navigation in the widely-used HM3D environment, as illustrated in Fig. 5. In this task, the agent needs to understand the detailed attributes of an object instance and navigate to an exact location.
Dataset Preparation.
Previous works on the instance-level object navigation task focus on discrete graph-based environments [64] and are limited by complicated human annotation process [11]. In this work, instead, we utilize the recent success of the Generative Vision-and-Language Models (GVLMs) [40, 35, 51] to automatically build the open-vocabulary text-goal setting for the InstanceNav task on the popular HM3D environment. Specifically, we randomly select a goal view for each validation episode to render an image of the goal object in the Instance Image Navigation (IIN) dataset [20]. Each episode in the IIN dataset corresponds to a unique goal object instance. To specify each object instance, we follow a previous attempt [38] to separate text descriptions into two aspects: Intrinsic Attributes and Extrinsic Attributes. Intrinsic attributes cover inherent characteristics of the object, such as shape, color, and material. Extrinsic attributes describe the environment surrounding the object, which is used to determine instances with similar intrinsic attributes. We instruct a GVLM model (i.e., CogVLM [51]) to generate both types of attributes according to the instance image with a hand-crafted prompt (See the Appendix for details). The original ground truth trajectories are preserved to construct the additional InstanceNav test set. In total, the test set of the text-goal setting comprises 1,000 episodes featuring 795 unique objects across 36 scenes.
Method with Mapping with LLM LLM Extra Sensors SR SPL
L3MVN [61] GPT-2 Depth, GPS 35.2 16.5
PixelNav [6] GPT-4 - 37.9 20.5
ESC [63] GPT-3.5 Depth, GPS 39.2 22.3
CoW [11] - Depth, GPS 6.1 3.9
ProcTHOR [9] - Depth, GPS 13.2 7.7
ZSON [37] - - 25.5 12.6
PSL (Ours) - - 42.4 19.2
Table 1: Comparison with state-of-the-art methods on the ObjectNav task. Our PSL surpasses both LLM-based and Mapping-based methods in terms of Success Rate(SR).
Results under the Text-goal Setting.
To demonstrate our PSL agent’s ability to understand complicated text descriptions and navigate to specified instances, we compare it with state-of-the-art zero-shot object-goal navigation methods under the text-goal setting of the InstanceNav task. In Table 2, we find that most of the existing zero-shot object navigation methods struggle in this task, including both map-based methods CoW [11] ($1.8\%$ in SR) and LLM-based method ESC [63] ($6.5\%$). All these methods perform inferior to the methods that receive semantic goals (i.e., ours and ZSON), indicating the effectiveness of end-to-end approaches that are pre-trained on large-scale unlabeled data in this task. Similar to our approach, we also transfer a well-trained ImageNav method (OVRL [57]) to perform the text-goal evaluation using our semantic expansion inference scheme. We find that though good at the ImageNav task, the OVRL model performs poorly when given retrieved image goals, indicating that the learnable goal encoder and observation encoder failed to encode semantic-level information for the InstanceNav task. In comparison, our PSL further brings a significant improvement over the ZSON baseline ($+5.9\%$ in SR) and achieves state-of-the-art results.
Method Backbone with LLM with Mapping Extra Sensors SR SPL
CoW [11] ViT-Base Depth, GPS 1.8 1.1
GoW [63] ViT-Base Depth, GPS 7.2 4.2
ESC [63] ViT-Base Depth, GPS 6.5 3.7
${}^{\dagger}$ [57] ResNet-50 - 3.7 1.8
ZSON [37] ResNet-50 - 10.6 4.9
PSL (Ours) ResNet-50 - 16.5 7.5
Table 2: Comparison with state-of-the-art methods in the text-goal navigation track of the InstanceNav task. We report the baseline results based on the released code and models. ${}^{\dagger}$e perform evaluation with our proposed semantic expansion inference scheme.
Results under the Image-goal Setting.
We then compare our PSL agent with other methods under the original Image-goal setting[20]. The agent is tasked to navigate to an object instance specified by an image. This task is different from the ImageNav task since the goal images are shot from perspectives with different camera configurations, including different pitch angles and HFOV. Existing state-of-the-art methods for this task heavily rely on supervision signals from scene object annotations and labeled viewpoints. For a fair comparison, we mainly focus on the performance of zero-shot approaches. Results in Table 3 show that our method brings significant improvements on state-of-the-art image-goal navigation methods ($+13.1\%$ compared to FGPrompt [55] and $+8.4\%$ compared to ZSON [37]). Meanwhile, we significantly reduce the gap with the supervised methods like OVRL-V2 in terms of SPL ($11.4\%$ vs. $11.8\%$), even though we use a much smaller vision backbone (ResNet-50 vs. ViT-Base).
Figure 5: Comparison of the InstanceNav task with the ObjectNav task. The ObjectNav task allows the agent to navigate to any chair in the scene, while the InstanceNav task has only one target object specified by detailed attribute descriptions.
Method Backbone Supervised Pre-training Data SR SPL
RL Agent [20] ResNet-18 - 8.3 3.5
OVRL-V2 [56] ViT-Base Gibson 24.8 11.8
OVRL-V2 [56] ViT-Base Gibson 0.6 0.2
${}^{\dagger}$ [57] ResNet-50 HM3D 8.0 4.2
FGPrompt [55] ResNet-9 HM3D 9.9 2.8
ZSON [37] ResNet-50 HM3D 14.6 7.3
PSL (Ours) ResNet-50 HM3D 23.0 11.4
Table 3: Comparison with state-of-the-art methods in the image-goal navigation track of the InstanceNav task. ${}^{\dagger}$e reimplement OVRL based on released pre-trained weight. In this track, the “Supervised” mark means human labels on the object and viewpoints are used.
5.4 Ablation Study
Effectiveness of Prioritized Semantic Learning.
To determine the effectiveness of our proposed method, we first conduct ablation studies under the image-goal setting of the InstanceNav task. In Table 4, we report the evaluation results of different variants of our method. We start from the ZSON [37] baseline, with a fixed CLIP goal encoder and a learnable ResNet-50 observation encoder. Then we apply the entropy-minimized goal view selection algorithm to select meaningful goal images during pre-training data pre-processing. This modification brings $+2.1\%$ improvement in success rate. We then add the perspective reward relaxation technique, which enhances goal orientation with an alternative pitch angle. This approach aims to prevent the agent from being immersed in exactly matching the goal view and observations, neglecting the semantic correlation between them. However, we discovered that with the vanilla model design, the augmented image goal confuses the agent and decreases the navigation success rate (from 12.7% to 10.8%). We argue that this phenomenon is due to the inferior agent design that neglects to reason semantic differences in the goal images and observation images. Incorporating with the PSL agent alleviates this issue. We found that it boosts the navigation success rate by $+3.8\%$ with the goal view selection approach, but more impressively, it cooperates well with the perspective relaxation mechanism and contributes a lot to the improvement ($+9.3\%$ in SR and $+4.2\%$ in SPL). We attribute it to the semantic perception module. This module captures and compares semantics in both goal and observation images with the same granularity, preventing the agent from learning sub-optimal navigation policies.
Method Input PSL Agent PST Strategy SR SPL
SPM GVS PRR
ZSON [37] Image 12.7 6.5
PSL (Ours) Image 14.8 7.7
Image 10.8 5.5
Image 16.5 6.5
Image 22.0 10.7
Table 4: Ablation studies of different components in our Prioritized Semantic Learning (PSL) agent and Prioritized Semantic Training (PST) strategy under the Image-Goal setting of the InstanceNav task. The default entry is marked in gray. “SPM”: Semantic Perception Module; “GVS”: Goal View Selection; “PRR”: Perspective Reward Selection.
Ablation on Semantic Expansion Inference.
To alleviate the granularity inconsistency issue of text and image embeddings during testing, we propose a semantic expansion inference scheme for our PSL agent. We perform elaborate ablation studies under the text-goal setting to investigate its effectiveness. We compare two different implementations of support set in the semantic expansion inference scheme, including collecting goal images from the image-goal navigation dataset, which refers to the ImageNav support set; and from the instance image navigation dataset, namely the IIN support set. In Table 5, we start with ablating the detailed configurations of the retrieval module and have two main findings: 1) using the support set images from the limited categories of objects as the dataset defined yields inferior results ($11.1\%$ in SR). Instead, the diversity of the ImageNav support set brings $+1.3\%$ improvements. 2) directly replacing the image features with text queries performs poorly ($6.6\%$ in SR) due to the modality gap [32]. Retrieving image features to augment the text queries yields better goal embeddings that share the same level of semantic granularity as training, thus significantly improving the navigation success rate by $+9.9\%$.
PSL Agent SEI Scheme Support Set #Supp. Vec. SR SPL
IIN 3.5K 11.1 5.4
ImageNav 0.1M 12.4 6.6
ImageNav 0.1M 6.6 2.7
ImageNav 0.1M 16.5 7.5
Table 5: Ablation studies of the Semantic Expansion Inference (SEI) scheme under the Text-Goal setting of the InstanceNav task.
5.5 Qualitative Analysis
Qualitative examples for PSL agent.
In Fig. 6 we present qualitative examples of our PSL agent navigating to two different object instances given detailed language instruction (e.g., , “Find the plant made of the black pot and grayish-white branches, around with a transparent glass dining table and four white chairs.”). The agent navigates across rooms, determining relative object instances according to the instruction, and finally stops near the object. We find that given detailed language instruction, the agent is able to differentiate between objects of the same category. For instance, in the first case, the agent bypasses a plant with “a white bottle and pink flower” which differs from the intrinsic attribute description. After a left turn near the first plant, the agent arrives at its destination and faces the correct plant instance with “black pot and grayish-white branches”. In another case, the agent hovers for a while in the intersection of bedroom and dining room to find the “white chair” and last the episode at “several white chairs with a glass dining table” that mentioned in the extrinsic attributes.
Figure 6: Qualitative examples for our PSL agent navigating to an object instance according to intrinsic and extrinsic object attributes in the InstanceNav dataset. For each trail, the agent is initialized at a random position in the room and given language instruction “Find the …”.
Figure 7: t-SNE vis. of the embeddings used by semantic expansion inference scheme.
Visualization of semantic expansion inference.
In Fig. 7, we provide t-SNE visualization of the embeddings used in our semantic expansion inference scheme. We present a triplet example where we retrieve an embedding from the support set using a text query. The retrieved embedding shows less difference from the ground-truth goal image embedding compared to the original text query. Our support set encompasses the object categories in InstanceNav and extends to a broader range of embeddings from various perspectives, contributing to the performance gain in navigation success rate.
6 Conclusion
We present PSL, a zero-shot approach for instance navigation. PSL is built upon the open-vocabulary and cross-modal ability of CLIP, and targets addressing the limited semantic perceiving ability of the observation encoder in existing methods. Our method is composed of three different parts, including a PSL agent with a semantic perception module module to reason the semantic differences between the observation and goal, a prioritized semantic training strategy with clear semantic supervision to train the PSL agent, and a semantic expansion inference scheme to maintain the granularity level of the goal-semantic. We investigate how this agent acts when instructed to navigate to a specific object by extending the zero-shot object navigation task to a zero-shot instance navigation task in the widely-used HM3D environment. we find that most existing methods struggle in this task, while our PSL agent outperforms other approaches.